# Let the Trial Begin: A Mock-Court Approach to Vulnerability Detection using LLM-Based Agents

### v1

@@ Abstract.
Detecting vulnerabilities in source code remains a critical yet challenging task, especially when benign and vulnerable functions share significant similarities. In this work, we introduce VulTrial, a courtroom-inspired multi-agent framework designed to enhance automated vulnerability detection. It employs four role-specific agents, which are security researcher, code author, moderator, and review board. Through extensive experiments using GPT-3.5 and GPT-4o we demonstrate that VulTrial outperforms single-agent and multi-agent baselines. Using GPT-4o, VulTrial improves the performance by 102.39% and 84.17% over its respective baseline. Additionally, we show that role-specific instruction tuning in multi-agent with small data (50 pair samples) improves the performance of VulTrial further by 139.89% and 118.30%. Furthermore, we analyze the impact of increasing the number of agent interactions on VulTrial’s overall performance. While multi-agent setups inherently incur higher costs due to increased token usage, our findings reveal that applying VulTrial to a cost-effective model like GPT-3.5 can improve its performance by 69.89% compared to GPT-4o in a single-agent setting, at a lower overall cost.
copyright: nonejournalyear: 2025doi: XXXXXXX.XXXXXXXconference: International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilisbn: 978-1-4503-XXXX-X/2018/06
@@ 1. Introduction
Vulnerability detection in software code is a critical task to ensure system security and prevent potential exploits (Chen et al., 2023). Manual vulnerability detection, however, can be time-consuming and error-prone, especially given the scale and complexity of modern software systems. To automate this process, researchers have explored various machine learning and deep learning approaches (Zou et al., 2021a; Russell et al., 2018; Chernis and Verma, 2018; Risse and Böhme, 2024; Zou et al., 2021b; Chakraborty et al., 2021). Recently, with substantial advancements in large language models (LLMs) and their impressive performance on natural language tasks (Wang et al., 2023; Abbasiantaeb et al., 2024; Zhuang et al., 2023; Neumann et al., 2024), several studies have attempted to utilize LLMs for automated vulnerability detection tasks (Tamberg and Bahsi, 2024; Ullah et al., 2023; Zhou et al., 2024; Nong et al., 2024; Zhang et al., 2024; Widyasari et al., 2024), which shows promising outcomes.
@@
Despite these advances, recent research by Ding et al. (Ding et al., 2025b) highlighted significant limitations when evaluating current LLM-based vulnerability detection approaches. Their study proposed a challenging new benchmark called PrimeVul, where the functional differences between benign and vulnerable code pairs are subtle, less than 20%. These cases where the code has a high degree of similarity are critical to evaluate because in real-world scenarios, vulnerabilities often stem from subtle, non-obvious code changes (Ni et al., 2023; Tan et al., 2024). On this dataset, LLM-based approaches exhibited low performance. The best performance comes from using GPT-4 (OpenAI, 2023b), which correctly predicts both codes in a pair (vulnerable and benign), referred to as Pair-wise Correct Prediction (P-C), 12.4% of cases. This LLM-based approach used in PrimeVul was implemented under the conventional single-agent system, wherein one model attempts to identify and explain vulnerabilities on its own.
Single-agent systems are inherently constrained by their reliance on a single perspective when analyzing vulnerabilities, which can lead to issues such as hallucination and overconfidence in incorrect reasoning (Zhang et al., 2023; Shi et al., 2025; Huang et al., 2025; Liang et al., 2023; Du et al., 2023). Moreover, these single-agent systems may fail to critically verify their own output, which is crucial when identifying subtle vulnerabilities (Liang et al., 2023).
Given these limitations, recent research has begun exploring multi-agent systems to enhance vulnerability detection (Wen et al., 2024; Hu et al., 2023; Wei et al., 2024). Multi-agent approaches distribute specific tasks among role-specific agents while allowing discussion and more information exchange between agents. GPTLens (Hu et al., 2023) is one such approach, dividing vulnerability detection into two stages, generation by an auditor agent and discrimination by a critic agent. In GPTLens, the final decision is made by the critic agent, which is tasked to be a harsh vulnerability critic and doubt the vulnerabilities and reasoning given by the auditor agent.
@@
In this work, we propose VulTrial, a novel multi-agent LLM framework that employs a courtroom-inspired scenario. VulTrial features four role-specific agents: the security researcher (prosecutor), code author (defense attorney), moderator (judge), and review board (jury). Unlike GPTLens, where the agent that does the critic delivers the final result, VulTrial delegates the final prediction to a dedicated review board agent, which is designed to be unbiased toward the security researcher and code author agents. The review board bases its decision on arguments from the security researcher (highlighting potential vulnerabilities), the code author (defending the code), and the moderator (summarizing debate points). Review board agent works similarly to the jury in the courtroom scenario, where they need to hear all the evidence presented in the case and make a final verdict based on it. Additionally, VulTrial allows for customizable rounds of interaction among the first three agents, which can provide opportunities for rebuttal and refinements.
We evaluate VulTrial results using two base LLMs, GPT-3.5 and GPT-4o (see Section 3.3). We benchmarked VulTrial against single-agent baselines from PrimeVul (Ding et al., 2025b) and a multi-agent baseline, GPTLens (Hu et al., 2023), all of which also utilize the same base LLMs for a fair comparison. PrimeVul (Ding et al., 2025b) (chain of thought and fine-tuning) and a multi-agent baseline, GPTLens (Hu et al., 2023). In the experiment using GPT-3.5, VulTrial outperforms the single-agent and multi-agent baselines by 151.69% and 239.78% respectively. Similarly, using GPT-4o, VulTrial also improves the performance of single-agent and multi-agent baselines by 102.39% and 84.17%.
Furthermore, to enhance VulTrial’s performance and analyze the impact of instruction tuning on a role-specific agent, we conducted additional experiments by applying instruction tuning to each agent in VulTrial. For the instruction tuning dataset, we selected 50 sample pairs from the validation data. The results indicate that tuning the moderator agent in VulTrial produced the most significant improvements, achieving an 18.53% increase compared to its non-tuned counterpart and outperforming the single-agent and multi-agent baselines by 139.89% and 118.30%, respectively. We also examined the effect of increasing the number of discussion rounds in VulTrial and observed that, when using GPT-4o, adding more rounds does not result in improved performance. Finally, an ablation study revealed that each agent within VulTrial is crucial for achieving the best overall performance. To summarize, the contributions of our study are as follows:
  • We propose VulTrial, a novel multi-agent, courtroom-inspired approach leveraging LLMs for automated vulnerability detection. VulTrial consists of role-specific agents (i.e., security researcher, code author, moderator, and review board). The ablation study highlights that each agent in VulTrial is necessary to get the best performance.
@@
  • We demonstrate that targeted instruction tuning of role-specific agents in VulTrial improves vulnerability detection performance compared to the baselines. Even with limited training examples (as few as 50 pairs), we get an improvement of 18.53% compared to the non-tuned version.
  • We conduct extensive experiments using GPT-3.5 and GPT-4o in PrimeVul dataset (Ding et al., 2025b) and compare the VulTrial with single-agent and multi-agent baselines. Overall, the results using VulTrial outperform the baselines.
@@ 2. VulTrial
Our proposed method, namely VulTrial, works by utilizing multiple agents, all of which are powered by LLMs. VulTrial consists of four role-specific agents, each inspired by distinct roles in a courtroom setting: a security researcher (prosecutor), code author (defense attorney), moderator (judge), and review board (jury). This structured debate encourages thorough, adversarial discussions, allowing for deeper analyses and refined conclusions regarding code vulnerabilities. Drawing inspiration from courtroom practices, where structured adversarial processes have been shown to reduce wrongful convictions and improve the fairness of judgments (Findley and Scott, 2006; Garrett, 2011), VulTrial aims to improve vulnerability detection in a similar manner. The overall architecture of these agents is illustrated in Fig. 1, with an example output in Fig. 2.
The first agent acts as a security researcher and is responsible to analyze the given code and identify potential vulnerabilities. The second agent, namely the code author agent, aims to provide counterarguments and justifications to refute the claims made by the security researcher agent. The third agent, which is the moderator agent, is tasked with summarizing the arguments presented by both the security researcher and the code author agents to ensure a structured exchange of viewpoints between them. VulTrial supports multiple rounds of interaction among the first three agents before passing their viewpoints into the last agent. The last agent, namely the review board agent, is tasked with analyzing all the information given by the moderator, the security researcher, and the code author agents. Based on this collective information, the review board delivers a final verdict regarding the security status of the code. This verdict highlights whether the potential vulnerability brought up by the security researcher agent is indeed a vulnerability. We chose to have the agents simulate a jury sentencing scenario over a judge sentencing scenario, as previous studies have shown that jury participation can help prevent excessively harsh punishments (Hans et al., 2015; McLeod, 2024). The details of each agent and their interaction will be discussed in the following subsections.
Figure 1. Overview of the VulTrial architecture. It employs a multi-agent framework inspired by a courtroom setting to assess code vulnerabilities. The security researcher (prosecutor) identifies potential vulnerabilities, the code author (defense attorney) provides counterarguments, the moderator (judge) summarizes key points, and the review board (jury) delivers the final verdict.
@@ 2.1. Security Researcher Agent (Prosecutor Role)
In VulTrial, the security researcher agent functions similarly to a prosecutor in a courtroom. A prosecutor’s role is to investigate evidence, establish the facts, and build a case against a defendant (i.e., the party that is accused in a court of law). In the case of vulnerability detection, the security researcher agent inspects the provided code, identifies potential vulnerabilities, and presents a rationale for why these vulnerabilities may pose a threat. An example of the security researcher agent’s output is shown in Fig. 2, where it highlights two potential vulnerabilities in the code, along with their reasoning and impact.
@@
The prompt used for this agent is highlighted in Fig. 3. In the first round, the agent only receives a code function and determines whether the code is vulnerable. In subsequent rounds, it also receives the output generated by two other agents (i.e., code author and moderator) which is highlighted as ¡chat history¿ in Fig. 3. Beyond simply reporting any vulnerability, the agent is explicitly asked to include (1) a description of each vulnerability, (2) the reasoning behind its identification, and (3) the potential impact if that vulnerability were to be exploited. We asked for this agent to include the reasoning as it has been highlighted in previous studies (Wei et al., 2022; Kojima et al., 2022; Cheng et al., 2024) that prompting LLMs to articulate their reasoning processes can significantly enhance their performance. Furthermore, including the potential impact of each vulnerability gives the other agents (i.e., the code author, moderator, and review board) a clearer perspective when they form their answer.
@@
You are the Security Researcher.
Identify all potential security vulnerabilities in the given code snippet.
Provide your output as a JSON array. Each element in the array represents one identified vulnerability and should include:
vulnerability”: A short name or description of the vulnerability. reason”: A detailed explanation of why this is a vulnerability and how it could be exploited. impact”: The potential consequences if this vulnerability were exploited. ¡code¿
¡chat history¿
Figure 3. Security researcher agent prompt.
2.2. Code Author Agent (Defense Attorney Role)
The code author agent in VulTrial acts similarly to a defense attorney in a courtroom scenario. While the prosecutor aims to prove that the defendant is guilty, the defense attorney counters these allegations by advocating for the defendant’s innocence. Applying this scenario to the vulnerability detection, the code author agent seeks to defend the integrity and security of their code, providing arguments and evidence to counter the security researcher agent’s claims. If the vulnerability is valid, the agent can acknowledge the issue and propose effective mitigation strategies.
The specific prompt for the code author agent is shown in Fig. 4. This prompt provides the agent with the code snippet under investigation along with information generated by the security researcher agent (included in ¡chat history¿), which outlines a list of potential vulnerabilities. The agent is then tasked with clearly articulating the following outputs: (1) identification of the vulnerability being discussed, (2) explicit acknowledgment or refutation of the vulnerability, and (3) detailed reasoning justifying the refutation or, in cases where the agent acknowledges the vulnerability, proposed methods for mitigation. An example of the code author agent’s output is also shown in Fig. 2, where it acknowledges one of the vulnerabilities identified by the security researcher and proposes a mitigation strategy. Additionally, it refutes the other vulnerability and provides reasoning for it.
Since the security researcher agent may highlight multiple potential vulnerabilities, it is crucial for the code author agent to clearly specify which vulnerability it addresses in each response. As emphasized by prior studies (Wei et al., 2022; Kojima et al., 2022; Cheng et al., 2024), explicitly prompting LLMs to articulate their reasoning has shown significant improvements in the accuracy and reliability of their outputs. Furthermore, clear explanations from the code author agent enable other agents (the moderator and review board) to better understand the justification or rationale behind the agent’s position. For the security researcher agent, this reasoning would be used for building their rebuttal response. This structured approach helps ensure that subsequent agents are equipped to offer well-informed counterarguments, summaries, or final evaluations.
You are the Code Author of the attached code.
The Security Researcher has presented a JSON array of alleged vulnerabilities. You must respond as if you are presenting your case to a group of decision-makers who will evaluate each claim. Your tone should be respectful, authoritative, and confident, as if you are defending the integrity of your work to a panel of experts.
For each identified vulnerability, produce a corresponding JSON object with the following fields:
vulnerability”: The same name/description from the Security Researcher’s entry. response-type”: ‘refutation’ if you believe this concern is unfounded, or ‘mitigation’ if you acknowledge it and propose a workable solution. reason”: A concise explanation of why the vulnerability is refuted or how you propose to mitigate it. ¡code¿
¡chat history¿
Figure 4. Code author agent prompt.
2.3. Moderator Agent (Judge Role)
In VulTrial, the moderator agent attempts to mirror the role of a judge in a courtroom scenario. Just as judges provide guidance to jurors and uphold fairness, within the context of vulnerability detection, this agent is tasked with objectively summarizing the arguments presented by both the security researcher agent (prosecutor) and the code author agent (defense attorney). It should focus only on the facts, avoiding any bias toward either side’s party. The primary purpose of this impartial summary is to assist the review board agent in reaching a decision, while also clarifying critical discussion points for subsequent rounds of argumentation by the security researcher and code author agents.
The prompt for this agent is highlighted in Fig. 5. In addition to accessing the code snippet in question, the moderator receives the security researcher’s identified vulnerabilities and the code author’s counterarguments (included in ¡code history¿). The moderator’s output consists of concise summaries of these two viewpoints, highlighting the key points from each. Fig. 2 presents an example of the moderator agent’s output. It summarizes the arguments from the security researcher (e.g., the number of potential vulnerabilities identified, their main reasons, and impacts) as well as from the code author, such as which vulnerabilities were acknowledged or refuted with the main reasoning.
You are the Moderator, and your role is to provide a neutral summary.
After reviewing both the Security Researcher’s identified vulnerabilities and the Code Author’s responses, provide a JSON object with two fields:
security researcher summary”: A concise summary of the vulnerabilities and reasoning presented by the Security Researcher. author summary”: A concise summary of the Code Author’s counterarguments or mitigation strategies. ¡code¿
¡chat history¿
Figure 5. Moderator agent prompt.

### v2

@@ Abstract.
Detecting vulnerabilities in source code remains a critical yet challenging task, especially when benign and vulnerable functions share significant similarities. In this work, we introduce VulTrial, a courtroom-inspired multi-agent framework designed to identify vulnerable code and to provide explanations. It employs four role-specific agents, which are security researcher, code author, moderator, and review board. Using GPT-4o as the base LLM, VulTrial almost doubles the efficacy of prior best-performing baselines. Additionally, we show that role-specific instruction tuning with small quantities of data significantly further boosts VulTrial’s efficacy. Our extensive experiments demonstrate the efficacy of VulTrial across different LLMs, including an open-source, in-house-deployable model (LLaMA-3.1-8B), as well as the high quality of its generated explanations and its ability to uncover multiple confirmed zero-day vulnerabilities in the wild.
vulnerability detection, large language models, multi-agent
journalyear: 2026copyright: ccconference: 2026 IEEE/ACM 48th International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilbooktitle: 2026 IEEE/ACM 48th International Conference on Software Engineering (ICSE ’26), April 12–18, 2026, Rio de Janeiro, Brazildoi: 10.1145/3744916.3773256isbn: 979-8-4007-2025-3/2026/04
@@ 1. Introduction
Vulnerability detection in software code is a critical task to ensure system security and prevent potential exploits (chen2023diversevul). Manual vulnerability detection can be time-consuming and error-prone, especially given the scale and complexity of modern software systems. To automate this, researchers have explored various machine learning and deep learning approaches (li2018vuldeepecker; russell2018automated; chernis2018machine; risse2024uncovering; zou2021interpreting; chakraborty2021deep). Recently, with substantial advancements in large language models (LLMs) and their impressive performance on natural language tasks (wang2023can; abbasiantaeb2024let; zhuang2023toolqa; neumann2024llm), several studies have found promising outcomes using LLMs for automated vulnerability detection (tamberg2024harnessing; ullah2023llms; zhou2024large_emerging; nong2024chain; zhang2024prompt; widyasari2024beyond)
@@
Despite these advances, recent research by Ding et al. (ding2024vulnerability) highlighted significant limitations when evaluating current LLM-based vulnerability detection approaches. Their study proposed a challenging new benchmark called PrimeVul, where the functional differences between benign and vulnerable code pairs are subtle, less than 20%. These cases where the code has a high degree of similarity are critical to evaluate because in real-world scenarios, vulnerabilities often stem from subtle, non-obvious code changes (ni2023distinguishing; tan2024similar). On this dataset, LLM-based approaches exhibited low performance. The best performance comes from using GPT-4 (openai2023gpt4), which achieves a Pair-wise Correct Prediction (P-C) score of 40 (out of a possible 435). P-C measures how many vulnerable-benign code pairs a method classifies entirely correctly, directly reflecting its ability to detect the subtle code changes that introduce a vulnerability. This LLM-based approach used in PrimeVul was implemented under the conventional single-agent system, wherein one model attempts to identify and explain vulnerabilities on its own. Single-agent systems are inherently constrained by their reliance on a single perspective when analyzing vulnerabilities, which can lead to issues such as hallucination and overconfidence in incorrect reasoning (zhang2023siren; shi2025mitigating; huang2025survey; liang2023encouraging; du2023improving). These single-agent systems may fail to critically verify their own output, which is crucial when identifying subtle vulnerabilities (liang2023encouraging).
@@
Given these limitations, recent research has begun exploring multi-agent systems to automate software engineering tasks (he2025llm; hu2023large; wei2024llm). Such approaches distribute tasks across role-specific agents while allowing discussion and richer information exchange between agents. In this work, we propose VulTrial, a novel multi-agent LLM framework that employs a courtroom-inspired scenario. VulTrial features four role-specific agents: the security researcher (prosecutor), code author (defense attorney), moderator (judge), and review board (jury). The review board bases its decision on arguments from the security researcher (highlighting potential vulnerabilities), the code author (defending the code), and the moderator (summarizing debate points). The review board agent works similarly to a jury in the courtroom scenario, where they need to hear all the evidence presented in the case and make a final verdict based on it. In VulTrial, we use the same underlying LLM for all agents, building on prior works (wei2022chain; chatterjee2024posix; hu2023large) that show prompt-induced roles can lead to different reasoning paths using the same model.
We benchmarked VulTrial against single-agent and multi-agent baselines. For the single-agent, we include deep learning (DL)-based models (CodeBERT (feng2020codebert), CodeT5 (wang2021codet5), UniXCoder (guo2022unixcoder), and LineVul (fu2022linevul)), as well as an LLM-based method (Ding et al.’s CoT prompts (ding2024vulnerability)). For the multi-agent setting, we utilized GPTLens (gptlens). All LLM-based methods are instantiated using GPT-3.5 and GPT-4o as the base model (see Section 3.3). The results show that VulTrial improves over the baselines. With GPT-4o, VulTrial nearly doubles the performance of both the best performing single-agent and multi-agent baselines (P-C score=81 vs. 40 and 44, respectively).
To assess the effect of instruction tuning on role-specific agents, we tuned each agent in VulTrial using 50 validation sample pairs. Instruction tuning the moderator agent in VulTrial improves the P-C score by 15 (18.52%) over the non-tuned version and surpasses the single-agent and multi-agent baselines by 56 (140%) and 52 (118%). To assess generalizability, we additionally confirmed the efficacy of VulTrial using LLaMA-3.1-8B, a publicly available model that can be deployed in-house. Separately, we perform an ablation study to verify that each agent contributes positively to achieving the best performance. Beyond quantitative results, we also qualitatively evaluate the generated explanations, examining completeness, clarity, actionability, and informativeness, and observe strong performance across all four dimensions. Finally, we deploy VulTrial in the wild, where it successfully uncovers multiple confirmed zero-day vulnerabilities, demonstrating its practical effectiveness.
To summarize, the contributions of our study are as follows:
  • We propose VulTrial, a novel multi-agent, courtroom-inspired approach leveraging LLMs for automated vulnerability detection. Also, different from many prior approaches that only output a binary label, VulTrial provides detailed explanations for each classification.
@@
  • We demonstrate that targeted instruction tuning of role-specific agents in VulTrial improves the performance.
  • We conduct extensive experiments demonstrating the efficacy of VulTrial on benchmark data and in the wild.
@@ 2. VulTrial
VulTrial leverages multiple LLM-based agents modeled after courtroom roles: a security researcher (prosecutor), code author (defense attorney), moderator (judge), and review board (jury). This structured debate enables thorough, adversarial discussions for deeper analyses and refined conclusions regarding code vulnerabilities. Drawing inspiration from courtroom practices, where structured adversarial processes have been shown to reduce wrongful convictions and improve the fairness of judgments (findley2006multiple; garrett2011convicting), VulTrial aims to improve vulnerability detection in a similar manner.
VulTrial architecture is shown in Fig. 1, with an example output in Fig. 2. The first agent acts as a security researcher and is responsible to analyze the given code and identify potential vulnerabilities. The second agent, namely the code author agent, aims to provide counterarguments and justifications to refute the claims made by the security researcher agent. The third agent, which is the moderator agent, is tasked with summarizing the arguments presented by both the security researcher and the code author agents to ensure a structured exchange of viewpoints between them. VulTrial supports multiple rounds of interaction among the first three agents before passing their viewpoints to the last agent. The last agent, namely the review board agent, is tasked with analyzing all the information given by the moderator, the security researcher, and the code author agents. Based on this collective information, the review board delivers a final verdict regarding the security status of the code. We chose to have the agents simulate a jury over a judge sentencing scenario, as previous studies have shown that jury participation can help prevent excessively harsh punishments (hans2015death; mcleod2024democratic). The details of the agent’s interaction will be discussed in the following subsections. The prompts can be found in our online appendix (ourreplicationpackageall).
Figure 1. VulTrial architecture which uses a courtroom-inspired multi-agent framework where the security researcher (prosecutor) identifies potential vulnerabilities, the code author (defense attorney) provides counterarguments, the moderator (judge) summarizes key points, and the review board (jury) delivers the final verdict on code vulnerabilities.
@@ 2.1. Security Researcher Agent (Prosecutor Role)
In VulTrial, the security researcher agent functions similarly to a prosecutor in a courtroom. A prosecutor’s role is to investigate evidence, establish the facts, and build a case against a defendant (i.e., the party that is accused in a court of law). In the case of vulnerability detection, the security researcher agent inspects the provided code, identifies potential vulnerabilities, and presents a rationale for why these vulnerabilities may pose a threat. An example is shown in Fig. 2, where it highlights two potential vulnerabilities in the code, along with their reasoning and impact.
@@
In the first round, the agent receives a code function and determines whether the code is vulnerable. In subsequent rounds, it also receives the output generated by two other agents (i.e., code author and moderator). Beyond simply reporting any vulnerability, the agent is explicitly asked to include (1) a description of each vulnerability, (2) the reasoning behind it, and (3) the potential impact if exploited. We require the agent to include its reasoning, as previous studies (wei2022chain; kojima2022large; cheng2024chainlm) show that prompting LLMs to articulate their reasoning can significantly enhance their performance. Furthermore, including the potential impact of each vulnerability gives the other agents a clearer perspective when they form their answer.
2.2. Code Author Agent (Defense Attorney Role)
The code author agent in VulTrial acts similarly to a defense attorney in a courtroom scenario. While the prosecutor aims to prove that the defendant is guilty, the defense attorney counters these allegations by advocating for the defendant’s innocence. Applying this scenario to vulnerability detection, the code author agent seeks to defend the integrity and security of their code, providing arguments and evidence to counter the security researcher agent’s claims. If the vulnerability is valid, the agent can acknowledge the issue and propose effective mitigation strategies.
The prompt provides the agent with the code snippet under investigation, along with information generated by the security researcher agent, which outlines a list of potential vulnerabilities (ourreplicationpackageall). The agent is then tasked with clearly articulating the following outputs: (1) identification of the vulnerability being discussed, (2) explicit acknowledgment or refutation of the vulnerability, and (3) detailed reasoning justifying the refutation or mitigation methods in case the vulnerability is acknowledged. An example of the code author agent’s output is also shown in Fig. 2, where it refutes one vulnerability and provides reasoning for it.
Since the security researcher agent may highlight multiple potential vulnerabilities, it is crucial for the code author agent to clearly specify which vulnerability it addresses in each response. Explanations from the code author agent enable other agents (the moderator and review board) to better understand the justification or rationale behind the agent’s position. For the security researcher agent, this reasoning would be used for building their rebuttal. This structured approach helps ensure that subsequent agents are equipped to offer well-informed counterarguments, summaries, or final evaluations.
2.3. Moderator Agent (Judge Role)
Just as judges provide guidance to jurors and uphold fairness in a courtroom scenario, within the context of vulnerability detection, this agent is tasked with objectively summarizing the arguments presented by both the security researcher agent (prosecutor) and the code author agent (defense attorney). It should focus only on the facts, avoiding any bias toward either side. This impartial summary helps the review board agent in reaching a decision, while also clarifying critical discussion points for subsequent rounds of argumentation by the security researcher and code author agents.
In addition to the code snippet, the moderator receives the security researcher’s identified vulnerabilities and the code author’s counterarguments. The moderator’s output consists of concise summaries of these two viewpoints, highlighting the key points from each. Fig. 2 shows an example of the moderator agent summarizing the arguments from the security researcher (e.g., the number of potential vulnerabilities identified, their main reasons, and impacts) as well as from the code author, such as which vulnerabilities were acknowledged or refuted with the main reasoning.